OpenStudy (anonymous):

Use the functions (x+1)/(x-3) and g(x)=(x+2)/(x+3) to find and simplify ex 1. (f+g)(x)

OpenStudy (anonymous):

Is f(x)=(x+1)/(x-3)

OpenStudy (anonymous):

yes, sorry about that

OpenStudy (anonymous):

Now, (f+g)(x)=f(x)+g(x)=(x+1)/(x-3) + (x+2)/(x+3) ={(x+1)(x+3) + (x+2)(x-3)}/{(x-3)(x+3)}

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

why over x-3 x+3?

OpenStudy (anonymous):

Take LCM

OpenStudy (anonymous):

is it (x+1)(x+2)?

OpenStudy (anonymous):

(f+g)(x)=f(x)+g(x)=(x+1)/(x-3) + (x+2)/(x+3) ={(x+1)(x+3) + (x+2)(x-3)}/{(x-3)(x+3)} ={x^2+3x+x+3+x^2-3x+2x-6}/{x^2-3^2} ={2x^2+3x-3}/{x^2-9}

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

nvm i got it

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

Yes thats the answer

OpenStudy (anonymous):

just simplified it

OpenStudy (anonymous):

@YesterdayiSaidTomorow

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

welcom